# Geolocation with Real Human Gameplay Data: A Large-Scale Dataset and Human-Like Reasoning Framework

### v2

@@
Abstract.
Geolocation, the task of identifying an image’s location, requires complex reasoning and is crucial for navigation, monitoring, and cultural preservation. However, current methods often produce coarse, imprecise, and non-interpretable localization. A major challenge lies in the quality and scale of existing geolocation datasets. These datasets are typically small-scale and automatically constructed, leading to noisy data and inconsistent task difficulty, with images that either reveal answers too easily or lack sufficient clues for reliable inference. To address these challenges, we introduce a comprehensive geolocation framework with three key components: GeoComp, a large-scale dataset; GeoCoT, a novel reasoning method; and GeoEval, an aspect-based metric designed to evaluate the correctness of the geolocation reasoning process. At the core of this framework is GeoComp (Geolocation Competition Dataset), a large-scale dataset collected from a geolocation game platform involving 740K users over two years. It comprises 25 million entries of metadata and 3.9 million geo-tagged locations spanning much of the globe, with each location annotated thousands to tens of thousands of times by human users. The dataset offers diverse difficulty levels for detailed analysis and highlights key gaps in current models. Building on this dataset, we propose Geographical Chain-of-Thought (GeoCoT), a multi-step reasoning framework designed to enhance the reasoning capabilities of Large Vision Models (LVMs) in geolocation tasks. GeoCoT improves performance by integrating contextual and spatial cues through a multi-step process that mimics human geolocation reasoning. Finally, we demonstrate that GeoCoT significantly boosts performance by up to 25% on classic geolocation metrics and by 9% in reasoning quality as measured by GeoEval:
Geolocation, Reasoning, Chain of thought, Dataset
copyright: noneisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Computer vision tasks
1. Introduction
@@
Google-WS-15k  (Clark et al., 2023a) 15k Global Map Service
GMCP (Zamir and Shah, 2014) 105K Local Map Service
StreetCLIP (Haas et al., 2023) 1M Unknown Map Service
Im2GPS  (Hays and Efros, 2008) 237 Local Web-Scraped
Im2GPS3K  (Vo et al., 2017b) 2997 Local Web-Scraped
YFCC4K  (Vo et al., 2017b) 4536 Local Web-Scraped
YFCC26K  (Theiner et al., 2022a) 26k Local Web-Scraped
MP-16  (Larson et al., 2017) 4.7M Local Web-Scraped
OSV-5M (Astruc et al., 2024) 5.1M Global Map Service
GeoComp 3.3M Global Map Service
@@
Table 1. Comparison of Existing Geolocation Datasets and GeoComp. “Local” refers to city- or region-specific data, while “Global” spans multiple continents. Darker green shades indicate broader geographic coverage.
Geolocation, the task of determining an image’s geographical location, is crucial for applications like crime tracking, navigation, fact-checking, and cultural exploration (Cheng et al., 2022; Chalvatzaras et al., 2022). It involves interpreting contextual clues within an image, such as architectural styles, road signs, natural landscapes, and cultural markers. Inferring location from such diverse indicators demands advanced reasoning, making geolocation a challenging task for both artificial models and human experts (Khan et al., 2024).
@@
Significant effort has been devoted to solving the geolocation task, but often at a coarse level of granularity. For example, methods like Im2GPS3K (Vo et al., 2017a) and PlaNet (Weyand et al., 2016a) frame the task by dividing the globe into grid cells and training deep neural networks to predict the correct cell for a given image. Subsequent studies improve precision by retrieving the most visually similar image from a dataset and using its coordinates as the predicted location (Zhu et al., 2022; Müller-Budack et al., 2018). The reason for this coarse granularity in many approaches is potentially due to the lack of high-quality datasets. For example, Im2GPS3K contains up to 35% non-localizable images (Astruc et al., 2024), while the YFC100M dataset includes irrelevant data such as indoor photos and food images, which provide little to no locational information (Theiner et al., 2022a). Additionally, many datasets are limited in size, with Georeasoner (Li et al., 2024a) featuring only 3K images, thereby restricting the robustness and generalizability of geolocation models. A comparison of these datasets is shown in Table 1.
@@
To address the above obstacles, in this work, we leverage the contributions of hundreds of thousands of geolocation game enthusiasts who provide real user prediction annotations while playing the game. Specifically, we launched a free, public-benefit-oriented online geoguessing platform in June 2022, as shown in Figure 1(a). A screenshot of the platform’s GUI is provided in Appendix A. In each game, two players independently guess the location based on the same image and their own hints, with scores determined by the distance between their predictions and the ground-truth location. The images are sourced from Google Maps, Baidu Maps, Tencent Maps, and Gaode Maps. The platform offers multiple game modes, allowing users to either choose opponents or join random matchups. As of December 17, 2024, this platform has 740,468 users, 3,954,397 locations as unique geolocation tasks, and 25,355,174 human response records. We name the collected dataset GeoComp. This rich and valuable dataset of real human responses enables us to evaluate task difficulty and filter out unreasonable cases. For instance, some tasks are too easy, such as when the name of a shopping mall in a city is clearly visible in the image, enabling most users to answer correctly. On the other hand, some tasks are highly challenging, where only a few users spend considerable time before providing accurate answers. Additionally, there are unreasonable tasks that contain no identifiable hints, making them unsolvable for all users despite significant effort.
@@
Our work makes key contributions to geolocation. First, we present GeoComp, a large-scale, human-annotated geolocation dataset with over 3.9 million location images with corresponding location labels, and 25 million human player annotations, featuring diverse geographic regions, languages, and environmental contexts. These annotations identify high-difficulty geolocation cases and establish benchmarks to guide future advancements. Second, we introduce the Geographical Chain of Thought (GeoCoT) framework, a multi-step reasoning approach that improves geolocation accuracy by leveraging geographical cues like landmarks, environmental features, and spatial relationships. Finally, through comprehensive evaluations involving human assessments and LLM inferences, we show that GeoCoT improves predictive performance by up to 25% while enhancing interpretability.
Figure 1. (a) The gaming logic of our platform: Two players independently guess the location based on the same image and their own hints, with scores determined by the distance between their predictions and the ground truth location. (b) The global map shows spatial heterogeneity, with dense clusters in more urbanized regions like Europe and Asia, and sparse coverage in areas like Africa and Siberia. (c) The pie chart highlights the proportional geo-tagged locations distribution, led by North America and Asia. (d) Unlike previous datasets like OSV-5M, where a single country (e.g., America) dominates 25% of the data, our dataset is balanced at country level.
2. Related work
2.1. Image Geolocation Task
@@
Image geolocation refers to determining the corresponding location of a given image, a crucial task in computer vision (Zhu et al., 2023b, a, 2024a), spatial data mining (Zhao et al., 2017, 2016; Zhang et al., 2023d; Han et al., 2023), and GeoAI (Zhao and Tang, 2017; Zhao et al., 2022; Zhang et al., 2023a, e). Previous research in image geolocalization could be primarily classified into two approaches: classification-based methods and retrieval-based methods. (1) Classification-based methods partition most regions of the Earth into multiple grid cells. Models are trained to classify each image into the correct cell (Clark et al., 2023b; Pramanick et al., 2022; Müller-Budack et al., 2018; Seo et al., 2018; Weyand et al., 2016b). The center coordinates of each cell are used as the predicted values. However, due to the limited number of cells, the granularity of the predicted values is coarse, preventing precise predictions. (2) Retrieval-based methods establish a database of geographic images with GPS coordinates. For a given input image, these methods retrieve the most similar image from the dataset and use its coordinates as the predicted location (Zhu et al., 2022; Müller-Budack et al., 2018; Zhang et al., 2023b; Workman et al., 2015; Liu and Li, 2019; Zhou et al., 2024). However, constructing a comprehensive global-level image database is clearly impractical.
2.2. Geolocation Dataset
@@
Existing geolocation datasets primarily originate from web-scraped or street-view images that have not been human-validated, raising concerns about their quality for effectively evaluating geolocation capabilities. For instance, datasets derived from web scraping, such as YFCC100M (Theiner et al., 2022b) and Im2GPS3K (Vo et al., 2017a), include a significant proportion of images depicting food, art, pets, and personal photographs. These types of images are often weakly localizable or entirely non-localizable (Theiner et al., 2022a). Street-view datasets also face limitations, such as restricted geographic coverage (Astruc et al., 2024); for example, (Mirowski et al., 2019)’s work includes data from only three cities in the United States. Furthermore, dataset collection processes often introduce biases. For instance, some commonly used platforms are inaccessible in certain countries, resulting in uneven geographic representation. Additionally, the difficulty of individual geolocation tasks varies widely within these datasets, but this aspect has not been comprehensively evaluated. For example, images taken at prominent landmarks are relatively easy to geolocate, while others offer no clear hints and are highly challenging (Astruc et al., 2024). These limitations undermine the reliability of current geolocation benchmarks.
Figure 2. Performance of game players of different levels in mainstream countries. Experts are defined as the top 15% in performance scores, while beginners are those in the bottom 15%.
2.3. Large Vision Language Models
@@
LLMs have exhibited extraordinary emergent abilities by scaling up data and model sizes, notably including instruction following (Li et al., 2023b; Dai et al., 2024), in-context learning (Brown et al., 2020), and Chain of Thought (CoT) (Kojima et al., 2022). Building on these emergent capabilities, significant research efforts have focused on enhancing cross-modality understanding and reasoning capabilities. Numerous studies have been conducted on various aspects of LVMs, encompassing structural design (Liu et al., 2024a; Cai et al., 2023; Liu et al., 2024c), data construction (LAION, 2023; Zhao et al., 2024), training strategies (Zhao et al., 2023; McKinzie et al., 2024; Lu et al., 2024), evaluations (Bithel and Bedathur, 2023), and the development of lightweight LVMs (Zhu et al., 2024b). Additionally, the robust capabilities of LVMs have been applied to other fields, such as medical image understanding (Zhou and Huang, 2024; Zhang et al., 2023c, 2024) and document parsing (Ye et al., 2023; Liu et al., 2024b). Furthermore, the development of multi-modal agents has advanced real-world applications, including embodied agents (Huang et al., 2023; Peng et al., 2024) and GUI agents (Yang et al., 2023; Li et al., 2024b; Song et al., 2024). However, the reasoning capabilities of LVMs in geolocation tasks remain underexplored. One of the primary reasons for this limitation is the lack of high-reasoning-value geographic data.
3. Data Overview
In this section, we first describe the data collection platform, then present the statistical distributions with visualizations, and finally showcase the performance of human players on the dataset, a unique feature that sets our dataset apart.
3.1. Geolocation Competition
@@
3.2. Dataset Statistic
Figure 1(a) presents a global heatmap of geo-tagged locations density, highlighting significant spatial heterogeneity in our dataset. High-density regions are concentrated in urbanized zones such as Europe, North America, and parts of Asia, whereas areas like Africa and Oceania are sparsely represented, often due to underdeveloped infrastructure or low population density. Figure 1(b) provides an overview of the proportional distribution of geo-tagged locations counts across continents, offering a macroscopic view of the dataset’s global spread. In Figure 1(c), we further illustrate the geo-tagged locations distribution by country. Notably, in contrast to datasets like OSV-5M (Astruc et al., 2024), which suffers from severe imbalances—such as the U.S. alone accounting for up to 25% of the total data—our dataset achieves a more balanced global distribution. No single country or continent dominates the dataset, ensuring a more equitable geographic representation and highlighting areas where further data collection efforts may be needed.
3.3. Human Player Performance
@@
Figure 3. Comparison of previous geolocation tasks and our proposed paradigm: while previous works focused on coarse-grained predictions limited by dataset quality, our generation and reasoning-based method enables fine-grained city-level predictions.
Our dataset not only includes image and location information but also rich human player performance data on the task. This label information serves not only as a valuable metric for evaluating the difficulty of different tasks but also as a benchmark for understanding human decision-making in geolocation challenges. In this subsection, we analyze the performance score across players and countries, providing insights into how human players perform on a global scale and how their accuracy varies across different regions and task types. We use GeoGuessr’s scoring formula to evaluate a user’s performance on a single task:
$S=\textstyle\left\lfloor\exp\left(-\frac{d}{s_{d}}\right)\times 5000\right\rfloor.$
Here, $S$ is the user’s score (0 to 5000), $d$ is the geographic distance between the predicted and actual locations (in kilometers), and $s_{d}$ is the maximum distance within the area divided by 10. The score decreases exponentially as $d$ increases. For example, $s_{d}$ is 1805 km globally (based on Earth’s maximum distance of 18,050 km) and 615 km for China, reflecting smaller scales. A perfect prediction ($d=0$) yields $S=5000$. A player’s performance score is defined as the average score across all their tasks. Similarly, a country’s performance score is the average score across all tasks performed within that country.
Player Performance Across Levels. The performance of game players across different levels, as illustrated in Figure 2, highlights significant gaps between beginners and experts in mainstream countries. Expert players, defined as the top 15% of performers, consistently achieve much higher performance metrics compared to beginners, defined as the bottom 15%, with noticeable gaps in countries like Canada, China, and India. For example, in Canada, experts perform nearly 10 times better than beginners, underscoring the steep learning curve involved in mastering the game. This performance gap presents challenges for new players, as it emphasizes the level of skill, strategy, and game knowledge required to compete effectively at higher levels.
Player Performance Across Countries. The player performance across countries, as shown in Figure 2, demonstrates significant variations influenced by three key factors: climate, geographic size, and language. Players tend to perform well in countries such as Germany, France, and Japan. These nations are characterized by unique languages and relatively small geographic sizes. The presence of distinctive languages on urban street signs provides clear linguistic clues, enabling players to quickly identify the country. Additionally, the compactness of these countries allows for more precise guesses, resulting in higher scores. In contrast, despite China’s unique language, its vast size and diverse climates make pinpointing specific locations challenging, leading to lower scores. Similarly, large countries like the USA, China, and Canada face additional challenges due to their shared temperate climates and extensive territories, where players often confuse them due to similar vegetation and climate, reducing accuracy.
Player Performance Across Tasks. From Figure 2, we can also observe significant variations in player performance across different tasks. In certain countries, player performance is relatively low, while in others, it is notably higher. This highlights the diversity in task difficulty within our dataset, offering valuable insights for assessing and categorizing task complexity.
Table 2. Comparison of Precision, Recall and F1 scores in country-level and city-level geolocation. The scores are represented as follows: best, second, and third. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).

### v3

@@
Abstract
Geolocation aims to identify an image’s location and requires complex reasoning, playing an important role in navigation, monitoring, and cultural preservation. However, existing methods often yield coarse and non-interpretable predictions. A key challenge is the limited quality and scale of current geolocation datasets, which are typically small, automatically constructed, and suffer from noise and inconsistent difficulty. To address these challenges, we introduce a comprehensive geolocation framework with three key components: GeoComp, a large-scale dataset; GeoCoT, a novel reasoning method; and GeoEval, designed to evaluate the correctness of the geolocation reasoning process. At the core of this framework is GeoComp, a large-scale dataset collected from a geolocation game platform involving 740K users over two years. It comprises 25 million entries of metadata and 2.7 million geo-tagged locations spanning much of the globe, with each location annotated thousands to tens of thousands of times by human users. Building on this dataset, we propose Geographical Chain-of-Thought (GeoCoT), a multi-step reasoning framework designed to enhance the reasoning capabilities of Large Vision Models (LVMs) in geolocation tasks. Finally, we demonstrate that GeoCoT significantly boosts performance by up to 25% on classic geolocation metrics and by 9% in reasoning quality as measured by GeoEval:  Github link.
Geolocation with Real Human Gameplay Data:
A Large-Scale Dataset and Human-Like Reasoning Framework
1 Introduction
Geolocation, the task of determining an image’s geographical location, is crucial for applications like crime tracking, navigation, fact-checking, and cultural exploration Cheng et al. (2022); Chalvatzaras et al. (2022). It involves interpreting contextual clues within an image, such as architectural styles, road signs, natural landscapes, and cultural markers. Inferring location from such diverse indicators demands advanced reasoning, making geolocation a challenging task for both artificial models and human experts Khan et al. (2024).
@@ 1 Introduction
Google-WS-15k  Clark et al. (2023a) 15k Global Map Service
GMCP Zamir and Shah (2014) 105K Local Map Service
StreetCLIP Haas et al. (2023) 1M Unknown Map Service
Im2GPS  Hays and Efros (2008) 237 Local Web-Scraped
Im2GPS3K  Vo et al. (2017b) 2997 Local Web-Scraped
YFCC4K  Vo et al. (2017b) 4536 Local Web-Scraped
YFCC26K  Theiner et al. (2022a) 26k Local Web-Scraped
MP-16  Larson et al. (2017) 4.7M Local Web-Scraped
OSV-5M Astruc et al. (2024) 5.1M Global Map Service
GeoComp 2.7M Global Map Service
@@
Table 1: Comparison of Existing Geolocation Datasets and GeoComp. “Local” refers to city- or region-specific data, while “Global” spans multiple continents. Darker green shades indicate broader geographic coverage.
@@
Significant effort has been devoted to solving the geolocation task Vo et al. (2017a); Weyand et al. (2016a); Zhu et al. (2022); Müller-Budack et al. (2018), but often at a coarse level of granularity and unlocalizability. The reason is potentially due to the lack of high-quality datasets. For example, Im2GPS3K contains up to 35% non-localizable images Astruc et al. (2024), while the YFC100M dataset includes irrelevant data such as indoor photos and food images, which provide little to no locational information Theiner et al. (2022a). Additionally, many datasets are limited in size, with Georeasoner Li et al. (2024) featuring only 3K images, thereby restricting the robustness and generalizability of geolocation models. A comparison of these datasets is shown in Table 1.
@@
To address the above obstacles, in this work, we leverage the contributions of hundreds of thousands of geolocation game enthusiasts who provide real user prediction annotations while playing the game. Specifically, we launched a free, public-benefit-oriented online geoguessing platform in June 2022. A screenshot of the platform’s GUI is provided in Appendix A. In each game, two players independently guess the location based on the same image and their own hints, with scores determined by the distance between their predictions and the ground-truth location. The images are sourced from Google Maps, Baidu Maps, Tencent Maps, and Gaode Maps. The platform offers multiple game modes, allowing users to either choose opponents or join random matchups. As of December 2024, this platform has 740K users, 2.7M locations as unique geolocation tasks, and 25M human response records. We name the collected dataset GeoComp. This rich and valuable dataset of real human responses enables us to evaluate task difficulty and filter out unreasonable cases. For instance, some tasks are too easy, such as when the name of a shopping mall in a city is clearly visible in the image, enabling most users to answer correctly. On the other hand, some tasks are highly challenging, where only a few users spend considerable time before providing accurate answers. Additionally, there are unreasonable tasks that contain no identifiable hints, making them unsolvable for all users despite significant effort.
@@
Our work makes key contributions to geolocation. First, we present GeoComp, a large-scale, human-annotated geolocation dataset with over 2.7 million location images with corresponding location labels, and 25 million human player annotations, featuring diverse geographic regions, languages, and environmental contexts. These annotations identify different difficulty geolocation cases and establish benchmarks to guide future advancements. Second, we introduce the Geographical Chain of Thought (GeoCoT) framework, a multi-step reasoning approach that improves geolocation accuracy by leveraging geographical cues like landmarks, environmental features, and spatial relationships. Finally, through comprehensive evaluations involving human assessments and LLM inferences, we show that GeoCoT improves predictive performance by up to 25% while enhancing interpretability.
Figure 1: Comparison of data distribution between our GeoComp dataset and the existing SOTA OSV-5M dataset. (a) Global geographical heatmaps illustrating image density. (b–c) Country-level statistical breakdowns for GeoComp and OSV-5M, respectively. The Geo-Pop Index, defined as the harmonic mean of Area Bias and Population Bias, is used to comprehensively assess distribution balance. Index values are color-coded to indicate balance levels: High Balance, Moderate Balance, Imbalance, and Critical Imbalance.
2 Related work
2.1 Image Geolocation Task
@@
Image geolocation refers to determining the corresponding location of a given image, a crucial task in computer vision Zhu et al. (2023b, a, 2024b), spatial data mining Zhao et al. (2017, 2016); Zhang et al. (2023d); Han et al. (2023), and GeoAI Zhao and Tang (2017); Zhao et al. (2022); Zhang et al. (2023c, e). Previous research in image geolocalization could be primarily classified into two approaches: classification-based methods and retrieval-based methods. (1) Classification-based methods partition most regions of the Earth into multiple grid cells. Models are trained to classify each image into the correct cell Clark et al. (2023b); Pramanick et al. (2022); Müller-Budack et al. (2018); Seo et al. (2018); Weyand et al. (2016b). The center coordinates of each cell are used as the predicted values. However, due to the limited number of cells, the granularity of the predicted values is coarse, preventing precise predictions. (2) Retrieval-based methods establish a database of geographic images with GPS coordinates. For a given input image, these methods retrieve the most similar image from the dataset and use its coordinates as the predicted location Zhu et al. (2022); Müller-Budack et al. (2018); Zhang et al. (2023a); Workman et al. (2015); Liu and Li (2019); Zhou et al. (2024). However, constructing a comprehensive global-level image database is clearly impractical.
2.2 Geolocation Dataset
@@
Existing geolocation datasets primarily originate from web-scraped or street-view images that have not been human-validated, raising concerns about their quality for effectively evaluating geolocation capabilities. For instance, datasets derived from web scraping, such as YFCC100M Theiner et al. (2022b) and Im2GPS3K Vo et al. (2017a), include a significant proportion of images depicting food, art, pets, and personal photographs. These types of images are often weakly localizable or entirely non-localizable Theiner et al. (2022a). Street-view datasets also face limitations, such as restricted geographic coverage Astruc et al. (2024); for example, Mirowski et al. (2019)’s work includes data from only three cities in the United States. Furthermore, dataset collection processes often introduce biases. For instance, some commonly used platforms are inaccessible in certain countries, resulting in uneven geographic representation. Additionally, the difficulty of individual geolocation tasks varies widely within these datasets, but this aspect has not been comprehensively evaluated. For example, images taken at prominent landmarks are relatively easy to geolocate, while others offer no clear hints and are highly challenging Astruc et al. (2024). These limitations undermine the reliability of current geolocation benchmarks.
2.3 Large Vision Language Models
@@
Numerous studies have been conducted on various aspects of LVMs, encompassing structural design Liu et al. (2024b); Cai et al. (2023); Liu et al. (2024a), data construction LAION (2023); Zhao et al. (2024), training strategies Zhao et al. (2023); McKinzie et al. (2024); Lu et al. (2024), evaluations Bithel and Bedathur (2023), and the development of lightweight LVMs Zhu et al. (2024a). Additionally, the robust capabilities of LVMs have been applied to other fields, such as medical image understanding Zhou and Huang (2024); Zhang et al. (2023b, 2024) and document parsing Ye et al. (2023); Liu et al. (2024c). However, the reasoning capabilities of LVMs in geolocation tasks remain underexplored. One of the primary reasons for this limitation is the lack of high-reasoning-value geographic data.
3 Data Overview
3.1 Geolocation Competition
@@
3.2 Dataset Statistics and Geographic Balance
To comprehensively evaluate the geographical distribution quality of our dataset, we conduct a comparative analysis with the existing SOTA dataset, OSV-5M Astruc et al. (2024). Figure 1(a) visualizes the global density of the dataset via geographical heatmaps. Compared to OSV-5M, which exhibits a skewed distribution towards Western regions, GeoComp demonstrates a significantly more comprehensive coverage. First, GeoComp successfully mitigates the data sparsity in under-represented areas, achieving dense coverage in East Asia and South America. Second, our improved global balance does not sacrifice the density of data in well-covered areas. As evidenced by the higher heatmap intensity, our dataset maintains superior sampling density even in Western regions (e.g., Europe and North America) compared to OSV-5M. This dual advantage, expanding global reach while intensifying local density, ensures that GeoComp is both spatially balanced and densely distributed.
To further validate the distribution quality beyond visual inspection, we propose the Area Bias and Population Bias ratios, calculated as the quotient of a country’s proportion in the dataset to its respective global share of land area or population. To synthesize spatial and population factors, we define the Geo-Pop Index as the harmonic mean of Area Bias and Population Bias, where 1.0 indicates perfect parity. We categorize distribution quality into four levels based on the magnitude of deviation:High Balance ($0.8\text{--}1.4$) denotes optimal alignment with real-world shares; Moderate Balance ($0.3\text{--}0.8$ or $1.4\text{--}3.0$) and Imbalance ($0.1\text{--}0.3$ or $3.0\text{--}4.5$) reflect increasing degrees of skew; while Critical Imbalance (<0.1<0.1 or >4.5>4.5) identifies extreme outliers that are either under-represented by over $10\times$ or over-sampled by more than $4.5\times$. The statistical breakdown reveals a severe polarization in OSV-5M. As shown in Figure 1(c), it exhibits Critical Imbalance in major nations: the United States is disproportionately over-represented (accounting for $24.90\%$ of the data with an Index of $4.74$), while China is severely under-represented (only $0.16\%$, Index $0.01$). In contrast, GeoComp effectively mitigates these imbalances, achieving a status of High Balance across the mainstream country (Figure 1(b)). For instance, the representation of China is restored to a reasonable $16.39\%$ (Index $1.38$), and the United States is normalized to $6.75\%$ (Index $1.31$), reflecting a distribution more aligned with real-world geography and population. Overall, the majority of countries in GeoComp fall within the High Balance range, demonstrating a stable distribution (Global Std. Dev. of $0.57$). Conversely, OSV-5M is predominantly characterized by Moderate Balance and Critical Imbalance, resulting in a significantly higher deviation of $1.37$.
@@ 3.2 Dataset Statistics and Geographic Balance
Figure 2: Comparison of previous geolocation tasks and our proposed paradigm: while previous works focused on coarse-grained predictions limited by dataset quality, our generation and reasoning-based method enables fine-grained city-level predictions.
3.3 Human Annotations
Distinct from previous works, GeoComp includes rich human gameplay data, serving as a vital benchmark for evaluating geolocation difficulty and image localizability. We analyze performance using the GeoGuessr scoring mechanism, where a user’s score $S$ decays exponentially with the distance error $d$. The score is calculated as $S=\lfloor\exp(-d/s_{d})\times 5000\rfloor$, where $s_{d}$ is a scale factor derived from the map’s maximum distance. A perfect prediction ($d=0$) yields 5,000 points. To establish a reliable difficulty benchmark, we visualized the dataset for the annotated locations and calculated the average score for each location. The resulting distribution is shown in Figure 3.
As illustrated in the histogram, the dataset exhibits a broad spectrum of difficulty levels. The 0–1,000 score range constitutes the largest portion of the data, accounting for approximately 27.6%. This high concentration of low scores reveals a substantial subset of hard samples, which are typically located in remote or rural areas and lack explicit semantic cues such as text signs. For these locations, accurate localization requires fine-grained reasoning over implicit environmental features, including vegetation patterns, soil texture, and architectural styles. The poor human performance in this range suggests that average players generally lack the specialized domain knowledge needed to interpret these subtle cues, which often results in near-random guessing. This characteristic highlights the opportunity for models to surpass human performance by effectively learning high-entropy environmental representations. Overall, this distribution confirms that GeoComp serves as a robust benchmark, providing sufficient challenging samples to stress state-of-the-art models while retaining a substantial volume of human-solvable data to support meaningful geographic feature learning.
Figure 3: Number of locations distributed across different average score intervals.